# Elastic Data Transfer Optimization with Hybrid Reinforcement Learning

### v2

@@ Abstract
Modern scientific data acquisition generates petabytes of data that must be transferred to geographically distant computing clusters. Conventional tools either rely on preconfigured sessions, which are difficult to tune for users without domain expertise, or they adaptively optimize only concurrency while ignoring other important parameters. We present LDM, an adaptive data transfer method that jointly considers multiple parameters. Our solution incorporates heuristic-based parallelism, infinite pipelining, and a deep reinforcement learning based concurrency optimizer. To make agent training practical, we introduce a lightweight network simulator that reduces training time to less than four minutes and provides a $2750\times$ speedup compared to online training. Experimental evaluation shows that LDMconsistently outperforms existing methods across diverse datasets, achieving up to $9.5\times$ higher throughput compared to state-of-the-art solutions.
@@ I Introduction
The modern scientific landscape is experiencing a data deluge, driven by advances in instrumentation, sensing technologies, and computational power that generate datasets at an unprecedented scale. This phenomenon, often referred to as “Big Data,” is commonly characterized by three dimensions: immense volume, high velocity of acquisition, and wide variety of data types. Research infrastructures such as high-energy physics experiments, large-scale genomics initiatives, and global sensor networks now routinely produce petabytes of data. For analysis, this data often must be transferred to geographically distant High-Performance Computing (HPC) clusters [ligo, ares, combustion, astronomy, des, lsst, atlas, belle2, ameriflux, esgf, esxsnmp]. Moving data at this scale requires extremely high network bandwidth. To address this need, Internet2 has already increased its backbone to 400 Gbps [internet2400Gbps], and ongoing efforts aim to extend network capacity to the terabit scale [esnet1000Gbps].
@@
To maximize the capabilities of modern transfer infrastructures, it is essential to carefully tune a variety of parameters, such as pipelining (transferring multiple files through a single connection) [TCP_Pipeline, farkas2002], parallelism (transferring one file in parts) [R_Lee01, R_Hacker05, R_Karrer06, R_Lu05] and concurrency (transferring multiple files at a time) [kosar04, Kosar09, R_Liu10]. These adjustments can be made entirely at the application layer without altering the underlying transfer protocols, and they play a crucial role in enhancing the overall efficiency and throughput of end-to-end data transfers. Most existing approaches to maximizing bandwidth utilization in high-speed networks rely solely on concurrency. This method can be effective for datasets with uniformly sized medium-to-large files. However, it performs poorly when the dataset contains a few very large files or a large number of small files, since it does not take parallelism or pipelining into account.
@@
The key challenge is to design a framework that remains user-friendly, since end users often lack the domain knowledge required to manually tune parameters for optimal performance. Another important consideration is that networks are not static. Bandwidth can fluctuate due to various factors, and a fixed parameter configuration may no longer remain optimal. For this reason, most recent works have modeled the problem as an online optimization task. Hasibul et al. [drl] addressed this problem using deep reinforcement learning. Although the approach is conceptually sound, it is impractical in practice because it requires day-long online training before the agent can be deployed.
@@
This paper presents LDM, a novel adaptive data transfer method that considers not only concurrency but also pipelining and parallelism. Parallelism is addressed using a heuristic approach based on file size, while pipelining is implemented as an infinite strategy. Concurrency is optimized through a policy-driven deep reinforcement learning (DRL) agent. To eliminate the impracticality of day-long online training [drl], we design a lightweight simulator that reduces training time to as little as four minutes.
@@
  • We demonstrate through extensive experiments that LDM outperforms existing methods across all dataset scenarios, achieving up to $9.5\times$ higher throughput compared to state-of-the-art solutions.
@@ II Related Work & Motivation
Stochastic approximation algorithms [rao2016experimental, yun2017data, esma-tcc, liu2018toward] offer promising alternatives to static solutions as they can discover optimal parameter configuration in the runtime. However, Existing solutions in this domain fall short to offer practical option due to long convergence time and failure to provide fairness and stability guarantees in shared environments. for example, Yun et al. proposed ProbData [yun2017data] to tune the number of parallel streams and buffer size for memory-to-memory TCP transfers using stochastic approximation. ProbData is able to explore the near-optimal configurations through sample transfers but it takes several hours to converge which makes it impractical to use as majority of the transfers in HPNs last less than few hours [liu2018cross]. Also, we have observed that background traffic changes drastically over several hours in shared networks [arslan2018high], so it may even fail to converge due to large variations in sample transfers.
Simplified Architecture Overwhelm Systems: The most recent studies also have implemented adaptive approaches using online black-box optimization algorithms such as Bayesian methods, Gradient Descent, and Deep Reinforcement Learning [arifuzzaman2023falcontpds, marlin, drl]. These methods offer high utilization with reduced overhead by periodically collecting socket statistics during transfers. The black-box design allows for instant deployment across various networks, as they don’t rely on historical data. These studies finds out that, Concurrency is the most impactful parameter for data transfers. As a result, these solutions narrow the problem to optimizing the number of Concurrency, which can offset suboptimal settings in other parameters. For instance, on a 100 Gbps link, an optimally tuned connection might achieve 10 Gbps, so the solution would require 10 parallel connections. But if TCP buffer size value is suboptimal, throughput per connection could drop to 1 Gbps, prompting the optimizer to increase concurrency to 100. The optimizer can very quickly, often as little in 30 seconds, find this target value due to reduced search space. While these solutions now can easily achieved high network utilization, the increasing concurrency values might significantly degrade the performance of the existing systems and I/O processes. Although extending these solutions to multi-parameter optimization is fairly straightforward, they often perform much worse because the optimizer spends significant time searching for the optimal configuration due to exponentially increasing search space.
@@ III Understanding The Parameters
Figure 1: Visual comparison of data transfer behaviors under different optimization strategies. Pipelining eliminates control-channel idle gaps, concurrency transfers multiple files simultaneously, and parallelism splits large files. LDM combines all three to maximize throughput.
@@ III-C Parallelism: Partitioning Large Files
IV LDM: Systems Design
We introduce LDM to address the throughput maximization problem in high-speed data transfer. It combines both high-level and low-level design strategies. At the high level, we introduce a novel pipeline that determines optimal, stable, and fast-converging transfer parameters. This control strategy is developed and refined using an offline simulator, which allows for efficient agent training. At the low level, we focus on the implementation of this strategy, incorporating robust mechanisms for handling pipelining and parallelism within the production system.
Figure 5: LDM optimizes concurrency, parallelism, and pipelining without being tuned by the end user.
IV-A Handling Transfer Parameters in LDM
In LDM, we address the three key transfer parameters—parallelism, pipelining, and concurrency—in unique ways. The overall design is illustrated in the flowchart shown in Figure 5.
@@ IV-A Handling Transfer Parameters in LDM
Pipelining. Pipelining determines how frequently connections are closed and re-established. In LDM, we use infinite pipelining: a connection remains open until the worker is stalled by the concurrency controller or the transfer is completed. This approach allows the congestion window to grow to its full potential, avoiding the penalty of restarting from the initial window each time a connection is closed.
Concurrency. Concurrency remains the central and most critical parameter, as both pipelining and parallelism are partially dependent on it. For example, parallelism becomes ineffective when concurrency is fixed at one, and pipelining suffers if concurrency is unstable and frequently turns connections on and off. In LDM, concurrency is dynamically determined by a PPO agent, which we describe in detail in the following subsection.
@@ IV-D PPO Agent Architecture
Policy-driven deep reinforcement learning (DRL) algorithms are well-suited for problems where the environment changes rapidly and a generalizable actor is required. Among these, we adopt Proximal Policy Optimization (PPO) [schulman2017ppo], one of the most widely used policy-based reinforcement learning algorithms due to its balance of stability and sample efficiency. Our PPO agent follows the standard PPO architecture, with the key components highlighted in this subsection.
@@ IV-D4 Value Network
Figure 6: LDM introduces offline training of a deep reinforcement learning agent to quickly learn the behavior of the real environment.
IV-E The LDM Workflow
After describing the low-level components, we now present the high-level system design of LDM, illustrated in Figure 6. The workflow consists of three phases: the exploration phase, the offline training phase, and the production phase.
@@ V Evaluation
Figure 7: Performance comparison of LDM, Falcon, DRL, and GD with Heuristic with 4096 files sized 256 MB in Fabric-testbed. Only concurrency methods like DRL can reach the top speed, but LDMuses 25% less concurrency.
@@
The goal of this evaluation is to demonstrate the effectiveness, robustness, and practicality of LDM across a wide range of real-world data transfer scenarios. We first quantify the end-to-end performance gains achieved by LDM on diverse datasets and network conditions. We then evaluate the generalizability of the system across different TCP congestion control algorithms, followed by an analysis of stability and fairness when multiple transfers share the network. To understand the contribution of each subsystem, we conduct a detailed study of pipelining, parallelism, and concurrency in isolation and in combination. Finally, we validate key design choices, including the use of a continuous state space for the reinforcement learning controller. Together, these results provide a comprehensive assessment of LDM and establish its suitability for high-performance data transfer environments. In the experiments described in this section, we used $K=1.02$, $B=10$, and chunksize=512MBchunksize=512~MB.
@@ V-A Experimental Setup
Our evaluation is carried out on two widely used research testbeds, summarized in Table I. The primary experiments are conducted on the Fabric testbed [fabric-2019] across two geographically distant sites: NEWY (New York, USA) and CERN (Geneva, Switzerland). Although the underlying Fabric links for these sites are provisioned at 100 Gbps and 50 Gbps respectively, Fabric operates as a shared experiment network where Layer 3 routed paths are shared among multiple experiments. As a result, the maximum stable throughputs observed end-to-end are lower than the physical link capacities, and the values reported in Table I reflect these empirically measured stable rates.
@@
To broaden the evaluation across different bandwidth-delay regimes, we also conduct experiments on a second Fabric path between MICH and MASS, which offers a stable throughput of 35 Gbps with a 33 ms RTT. Additionally, to assess LDM under constrained network conditions, we include experiments on CloudLab between Clemson and Utah, which provides a 1 Gbps link with a 67 ms RTT. Together, these paths span a wide range of bandwidth-delay products, enabling a comprehensive assessment of LDM across both high-performance and limited-bandwidth environments.
@@ V-B Training in the Simulated Environment
A (Uniform) 1 TB
4096 $\times$ 256 MB
B (Uniform and Large) 1 TB
804 $\times$ 1 GB
4 $\times$ 55 GB
C (Small) 128 GB
238,770 $\times$ {100 KB – 1 MB}
D (Mixed) 256 GB
2 $\times$ 12 GB
3872 $\times$ {10 MB – 100 MB}
23,624 $\times$ {100 KB – 1 MB}
@@ V-C End-to-End Performance on Diverse Datasets
We evaluated LDM’s performance in terms of throughput maximization, early convergence, and stability across multiple emulated dataset scenarios. For benchmarking, we compared our approach against Falcon [arifuzzaman2023falcontpds], a recent state-of-the-art solution; Globus [globusonline], the most widely used transfer service; the DRL-based method by Hasibul et al. [drl]; and a gradient descent optimization strategy with heuristics enabled. In this section, we will refer to Hasibul’s method as DRL and the gradient descent optimization strategy with heuristics enabled as GD (Heuristic). Among the evaluated methods, only Globus operates in a static manner, whereas all others are adaptive. To evaluate the datasets using Globus, we initially attempted benchmarking with Globus Online. However, even after disabling verification, the transfer speed remained very slow. Therefore, for our experiments, we adopted the globus-url-copy utility from the community-maintained open-source Grid Community Toolkit (GCT 6.2). As discussed in Section II, we did not include FDT or MDTM in our comparisons due to the unavailability of their software.
We evaluated LDMusing the two Fabric testbeds (NEWY–CERN and MICH–MASS) described earlier to capture performance across high-latency and moderate-latency wide-area links. We used BBR as the primary congestion control algorithm because it is specifically designed to operate efficiently in high-bandwidth environments that mirror our target use cases.
@@ V-C1 Effectiveness of Agent-Based Optimization
Among adaptive methods, DRL achieves nearly the same throughput as LDM, but it does so using $\approx$ 33% more concurrency in both the testbeds, thereby validating our initial hypothesis. In the MICH–MASS path, DRL performs better than LDM, and this difference can be explained using the bandwidth–delay product (BDP). On the NEWY–CERN path, the maximum achievable congestion window for a single stream is approximately 3000 Mbps (throttled speed) multiplied by 85 ms RTT, which equals 255 Mb, or about 32 MB. On the MICH–MASS path, the corresponding value is 1500 Mbps multiplied by 33 ms RTT, which equals 49.5 Mb, or about 6 MB. In the first case, the file size is roughly eight times larger than the maximum congestion window, while in the second case it is more than forty times larger. As a result, on the MICH–MASS path, the slow-start overhead becomes negligible compared to the additional processing overhead introduced by infinite pipelining and parallelism, making DRL appear faster. However, such homogeneous file distributions are uncommon in real scientific workloads, and the advantage does not generalize beyond this specific scenario.
Falcon and GD (Heuristic) perform worse because gradient descent-based online optimization requires additional time to identify the concurrency value that maximizes throughput. As demonstrated in Figure 7, Falcon and GD (Heuristic) only reached their mean throughput after 46 seconds and 40 seconds, respectively, compared to just 3 seconds and 6 seconds for DRL and LDM. This figure illustrates the best runs of each method of the NEWY-CERN path in terms of throughput.
Finally, the standard deviation of concurrency further illustrates the effectiveness of agent-based methods. The standard deviations observed for PPO-based methods (DRL and LDM) are less than half of those for their gradient descent counterparts, demonstrating improved stability in optimization.

### v3

@@ Abstract
Modern scientific data acquisition generates petabytes of data that must be transferred to geographically distant computing clusters. Conventional tools either rely on preconfigured sessions, which are difficult to tune for users without domain expertise, or they adaptively optimize only concurrency while ignoring other important parameters. We present EDT, an adaptive data transfer method that jointly considers multiple parameters. Our solution incorporates heuristic-based parallelism, infinite pipelining, and a deep reinforcement learning based concurrency optimizer. To make agent training practical, we introduce a lightweight network simulator that reduces training time to less than four minutes and provides a $2750\times$ speedup compared to online training. Experimental evaluation shows that EDT consistently outperforms existing methods across diverse datasets, achieving up to $9.5\times$ higher throughput compared to state-of-the-art solutions.
@@ I Introduction
The modern scientific landscape is experiencing a data deluge, driven by advances in instrumentation, sensing technologies, and computational power that generate datasets at an unprecedented scale. This phenomenon, often referred to as “Big Data,” is commonly characterized by three dimensions: immense volume, high velocity of acquisition, and wide variety of data types. Research infrastructures such as high-energy physics experiments, large-scale genomics initiatives, and global sensor networks now routinely produce petabytes of data. For analysis, this data often must be transferred to geographically distant High-Performance Computing (HPC) clusters [33, 23, 34, 21, 29, 31, 27, 28, 26, 32, 30]. Moving data at this scale requires extremely high network bandwidth. To address this need, Internet2 has already increased its backbone to 400 Gbps [12], and ongoing efforts aim to extend network capacity to the terabit scale [8].
@@
To maximize the capabilities of modern transfer infrastructures, it is essential to carefully tune a variety of parameters, such as pipelining (transferring multiple files through a single connection) [10, 9], parallelism (transferring one file in parts) [17, 11, 14, 22] and concurrency (transferring multiple files at a time) [16, 15, 18]. These adjustments can be made entirely at the application layer without altering the underlying transfer protocols, and they play a crucial role in enhancing the overall efficiency and throughput of end-to-end data transfers. Most existing approaches to maximizing bandwidth utilization in high-speed networks rely solely on concurrency. This method can be effective for datasets with uniformly sized medium-to-large files. However, it performs poorly when the dataset contains a few very large files or a large number of small files, since it does not take parallelism or pipelining into account.
@@
The key challenge is to design a framework that remains user-friendly, since end users often lack the domain knowledge required to manually tune parameters for optimal performance. Another important consideration is that networks are not static. Bandwidth can fluctuate due to various factors, and a fixed parameter configuration may no longer remain optimal. For this reason, most recent works have modeled the problem as an online optimization task. Hasibul et al. [13] addressed this problem using deep reinforcement learning. Although the approach is conceptually sound, it is impractical in practice because it requires day-long online training before the agent can be deployed.
@@
This paper presents EDT, a novel adaptive data transfer method that considers not only concurrency but also pipelining and parallelism. Parallelism is addressed using a heuristic approach based on file size, while pipelining is implemented as an infinite strategy. Concurrency is optimized through a policy-driven deep reinforcement learning (DRL) agent. To eliminate the impracticality of day-long online training [13], we design a lightweight simulator that reduces training time to as little as four minutes.
@@
  • We demonstrate through extensive experiments that EDT outperforms existing methods across all dataset scenarios, achieving up to $9.5\times$ higher throughput compared to state-of-the-art solutions.
@@ II Related Work & Motivation
Stochastic approximation algorithms [24, 36, 35, 19] offer promising alternatives to static solutions as they can discover optimal parameter configuration in the runtime. However, Existing solutions in this domain fall short to offer practical option due to long convergence time and failure to provide fairness and stability guarantees in shared environments. for example, Yun et al. proposed ProbData [36] to tune the number of parallel streams and buffer size for memory-to-memory TCP transfers using stochastic approximation. ProbData is able to explore the near-optimal configurations through sample transfers but it takes several hours to converge which makes it impractical to use as majority of the transfers in HPNs last less than few hours [20]. Also, we have observed that background traffic changes drastically over several hours in shared networks [6], so it may even fail to converge due to large variations in sample transfers.
Simplified Architecture Overwhelm Systems: The most recent studies also have implemented adaptive approaches using online black-box optimization algorithms such as Bayesian methods, Gradient Descent, and Deep Reinforcement Learning [5, 4, 13, 2]. These methods offer high utilization with reduced overhead by periodically collecting socket statistics during transfers. The black-box design allows for instant deployment across various networks, as they don’t rely on historical data. These studies finds out that, Concurrency is the most impactful parameter for data transfers. As a result, these solutions narrow the problem to optimizing the number of Concurrency, which can offset suboptimal settings in other parameters. For instance, on a 100 Gbps link, an optimally tuned connection might achieve 10 Gbps, so the solution would require 10 parallel connections. But if TCP buffer size value is suboptimal, throughput per connection could drop to 1 Gbps, prompting the optimizer to increase concurrency to 100. The optimizer can very quickly, often as little in 30 seconds, find this target value due to reduced search space. While these solutions now can easily achieved high network utilization, the increasing concurrency values might significantly degrade the performance of the existing systems and I/O processes. Although extending these solutions to multi-parameter optimization is fairly straightforward, they often perform much worse because the optimizer spends significant time searching for the optimal configuration due to exponentially increasing search space.
@@ III Understanding The Parameters
Figure 1: Visual comparison of data transfer behaviors under different optimization strategies. Pipelining eliminates control-channel idle gaps, concurrency transfers multiple files simultaneously, and parallelism splits large files. EDT combines all three to maximize throughput.
@@ III-C Parallelism: Partitioning Large Files
IV EDT: Systems Design
We introduce EDT to address the throughput maximization problem in high-speed data transfer. It combines both high-level and low-level design strategies. At the high level, we introduce a novel pipeline that determines optimal, stable, and fast-converging transfer parameters. This control strategy is developed and refined using an offline simulator, which allows for efficient agent training. At the low level, we focus on the implementation of this strategy, incorporating robust mechanisms for handling pipelining and parallelism within the production system.
Figure 5: EDT optimizes concurrency, parallelism, and pipelining without being tuned by the end user.
IV-A Handling Transfer Parameters in EDT
In EDT, we address the three key transfer parameters—parallelism, pipelining, and concurrency—in unique ways. The overall design is illustrated in the flowchart shown in Figure 5.
@@ IV-A Handling Transfer Parameters in EDT
Pipelining. Pipelining determines how frequently connections are closed and re-established. In EDT, we use infinite pipelining: a connection remains open until the worker is stalled by the concurrency controller or the transfer is completed. This approach allows the congestion window to grow to its full potential, avoiding the penalty of restarting from the initial window each time a connection is closed.
Concurrency. Concurrency remains the central and most critical parameter, as both pipelining and parallelism are partially dependent on it. For example, parallelism becomes ineffective when concurrency is fixed at one, and pipelining suffers if concurrency is unstable and frequently turns connections on and off. In EDT, concurrency is dynamically determined by a PPO agent, which we describe in detail in the following subsection.
@@ IV-D PPO Agent Architecture
Policy-driven deep reinforcement learning (DRL) algorithms are well-suited for problems where the environment changes rapidly and a generalizable actor is required. Among these, we adopt Proximal Policy Optimization (PPO) [25], one of the most widely used policy-based reinforcement learning algorithms due to its balance of stability and sample efficiency. Our PPO agent follows the standard PPO architecture, with the key components highlighted in this subsection.
@@ IV-D4 Value Network
Figure 6: EDT introduces offline training of a deep reinforcement learning agent to quickly learn the behavior of the real environment.
IV-E The EDT Workflow
After describing the low-level components, we now present the high-level system design of EDT, illustrated in Figure 6. The workflow consists of three phases: the exploration phase, the offline training phase, and the production phase.
@@ V Evaluation
Figure 7: Performance comparison of EDT, Falcon, DRL, and GD with Heuristic with 4096 files sized 256 MB in Fabric-testbed. Only concurrency methods like DRL can reach the top speed, but EDTuses 25% less concurrency.
@@
The goal of this evaluation is to demonstrate the effectiveness, robustness, and practicality of EDT across a wide range of real-world data transfer scenarios. We first quantify the end-to-end performance gains achieved by EDT on diverse datasets and network conditions. We then evaluate the generalizability of the system across different TCP congestion control algorithms, followed by an analysis of stability and fairness when multiple transfers share the network. To understand the contribution of each subsystem, we conduct a detailed study of pipelining, parallelism, and concurrency in isolation and in combination. Finally, we validate key design choices, including the use of a continuous state space for the reinforcement learning controller. Together, these results provide a comprehensive assessment of EDT and establish its suitability for high-performance data transfer environments. In the experiments described in this section, we used $K=1.02$, $B=10$, and chunksize=512MBchunksize=512~MB.
@@ V-A Experimental Setup
Our evaluation is carried out on two widely used research testbeds, summarized in Table I. The primary experiments are conducted on the Fabric testbed [7] across two geographically distant sites: NEWY (New York, USA) and CERN (Geneva, Switzerland). Although the underlying Fabric links for these sites are provisioned at 100 Gbps and 50 Gbps respectively, Fabric operates as a shared experiment network where Layer 3 routed paths are shared among multiple experiments. As a result, the maximum stable throughputs observed end-to-end are lower than the physical link capacities, and the values reported in Table I reflect these empirically measured stable rates.
@@
To broaden the evaluation across different bandwidth-delay regimes, we also conduct experiments on a second Fabric path between MICH and MASS, which offers a stable throughput of 35 Gbps with a 33 ms RTT. Additionally, to assess EDT under constrained network conditions, we include experiments on CloudLab between Clemson and Utah, which provides a 1 Gbps link with a 67 ms RTT. Together, these paths span a wide range of bandwidth-delay products, enabling a comprehensive assessment of EDT across both high-performance and limited-bandwidth environments.
@@ V-B Training in the Simulated Environment
A (Uniform) 1 TB 4096 $\times$ 256 MB
B (Uniform and Large) 1 TB 804 $\times$ 1 GB 4 $\times$ 55 GB
C (Small) 128 GB 238,770 $\times$ {100 KB – 1 MB}
D (Mixed) 256 GB 2 $\times$ 12 GB 3872 $\times$ {10 MB – 100 MB} 23,624 $\times$ {100 KB – 1 MB}
@@ V-C End-to-End Performance on Diverse Datasets
We evaluated EDT’s performance in terms of throughput maximization, early convergence, and stability across multiple emulated dataset scenarios. For benchmarking, we compared our approach against Falcon [5], a recent state-of-the-art solution; Globus [1], the most widely used transfer service; the DRL-based method by Hasibul et al. [13]; and a gradient descent optimization strategy with heuristics enabled. In this section, we will refer to Hasibul’s method as DRL and the gradient descent optimization strategy with heuristics enabled as GD (Heuristic). Among the evaluated methods, only Globus operates in a static manner, whereas all others are adaptive. To evaluate the datasets using Globus, we initially attempted benchmarking with Globus Online. However, even after disabling verification, the transfer speed remained very slow. Therefore, for our experiments, we adopted the globus-url-copy utility from the community-maintained open-source Grid Community Toolkit (GCT 6.2). As discussed in Section II, we did not include FDT or MDTM in our comparisons due to the unavailability of their software.
We evaluated EDTusing the two Fabric testbeds (NEWY–CERN and MICH–MASS) described earlier to capture performance across high-latency and moderate-latency wide-area links. We used BBR as the primary congestion control algorithm because it is specifically designed to operate efficiently in high-bandwidth environments that mirror our target use cases.
@@ V-C1 Effectiveness of Agent-Based Optimization
Among adaptive methods, DRL achieves nearly the same throughput as EDT, but it does so using $\approx$ 33% more concurrency in both the testbeds, thereby validating our initial hypothesis. In the MICH–MASS path, DRL performs better than EDT, and this difference can be explained using the bandwidth–delay product (BDP). On the NEWY–CERN path, the maximum achievable congestion window for a single stream is approximately 3000 Mbps (throttled speed) multiplied by 85 ms RTT, which equals 255 Mb, or about 32 MB. On the MICH–MASS path, the corresponding value is 1500 Mbps multiplied by 33 ms RTT, which equals 49.5 Mb, or about 6 MB. In the first case, the file size is roughly eight times larger than the maximum congestion window, while in the second case it is more than forty times larger. As a result, on the MICH–MASS path, the slow-start overhead becomes negligible compared to the additional processing overhead introduced by infinite pipelining and parallelism, making DRL appear faster. However, such homogeneous file distributions are uncommon in real scientific workloads, and the advantage does not generalize beyond this specific scenario.
Falcon and GD (Heuristic) perform worse because gradient descent-based online optimization requires additional time to identify the concurrency value that maximizes throughput. As demonstrated in Figure 7, Falcon and GD (Heuristic) only reached their mean throughput after 46 seconds and 40 seconds, respectively, compared to just 3 seconds and 6 seconds for DRL and EDT. This figure illustrates the best runs of each method of the NEWY-CERN path in terms of throughput.
Finally, the standard deviation of concurrency further illustrates the effectiveness of agent-based methods. The standard deviations observed for PPO-based methods (DRL and EDT) are less than half of those for their gradient descent counterparts, demonstrating improved stability in optimization.